# Why Are Linear RNNs More Parallelizable?

### 04 Mar 2026 v1

@@ 2.3 Linear RNNs
Definition 4 leaves open the specific parameterization of AA and $b$ as a function of $\mathbf{x}_{t}$. The choice of parameterization is crucial for theoretical expressivity (Merrill et al., 2024) as well practical concerns like memory. For instance, constrained parameterizations (time-invariant or diagonal) were explored in early LRNNs, but these were shown to limit expressivity to $\mathsf{TC}^{0}$ (Merrill et al., 2024). Since then, newer LRNNs have adopted richer parameterizations that are non-diagonal but still efficient. Two broad classes of “slightly non-diagonal” parameterizations are diagonal-plus-low-rank (DPLR) and permutation-diagonal (PD) LRNNS. [^2] Henceforth, for brevity, we will write $\mathbf{A}_{t}$, $\mathbf{k}_{t}$, etc. to mean $A_{t}(\mathbf{x}_{t})$, $k_{t}(\mathbf{x}_{t})$, etc., leaving the dependence on $\mathbf{x}_{t}$ implicit.
@@ 5.1 DPLR LRNNs are $\mathsf{PNC}^{1}$-Complete
The input is a token stream of length $9n$ over $\{-1,0,1\}$, partitioned into $n$ consecutive blocks of length 99. Block $t$ encodes a matrix $\mathbf{A}_{t}\in\mathbb{K}^{3\times 3}$. The iterated matrix multiplication problem is to compute $\mathbf{P}_{n}=\prod_{t=1}^{n}\mathbf{A}_{t}\in\mathbb{Z}^{3\times 3}$.
@@
There exist a 44-layer RWKV-7 and a 44-layer DeltaNet that solves the iterated $3\times 3$ matrix multiplication problem and, in particular, checks if the $(0,0)$ entry of $\mathbf{P}_{n}$ is positive.
@@
For any $n$-state WFA $\mathcal{A}$ over $\mathbb{Q}$, there exist a 44-layer RWKV-7 and a 4-layer DeltaNet that both compute $f_{\mathcal{A}}(w)$.
@@ 6 Experiments
To better evaluate learning capacity, we construct balanced datasets with length generalization. The train set contains 70K samples uniformly drawn from $N\in[1,100]$ and the validation set contains 20K samples drawn from the same range. We evaluated on three test splits, each containing 10K samples, drawn from $[1,100]$, $[101,200]$, and $[201,300]$. Here, NN denotes the number of nodes in the graph task and the size of the matrix in the matrix multiplication task.
@@ A.2 Logarithmic Precision
The above reduction can easily be seen to be implementable by a logspace-uniform $\mathsf{AC}^{0}$ reduction. The values hnhn (for each $h=1,\ldots,m$) can be precomputed by a logspace machine since this depends only on the input size (more precisely, one can enumerate enough of hnhn for up to $m=O(n)$). The $\mathsf{AC}^{0}$ circuit simply needs to perform simple additions. ∎
@@
At the following separator it tests whether $S=0$ using the zero-mask. If $S=0$ (i.e. $s_{m-1}=i_{m}$), it moves to a state $q_{j}^{=}$ and on each symbol of the $j_{m}$-block simply does $S\leftarrow S+1$ (leaving II unchanged); since $S$ starts at 0, after $0^{j_{m}}$ we have $S=j_{m}$, and at the next separator it resets $I\leftarrow 0$ and returns to $q_{\mathsf{edge}}$, so $s_{m}=j_{m}$. If instead $S\neq 0$, the machine moves to a state $q_{j}^{\neq}$. On each symbol of the $j_{m}$-block it now performs
@@
where the choice again uses only the zero-mask bit of II. Initially $I=i_{m}$, so after $\min\{i_{m},j_{m}\}$ symbols we have added that many units back to $S$ and subtracted them from II. Because $j_{m}\geq i_{m}$, by the end of the $j_{m}$-block we have $I=0$ and
@@ C.1 Preliminaries
The input is a token stream of length $9N$ over alphabet $\mathbb{Z}$, partitioned into NN consecutive blocks of length 99. The $t$-th block encodes a matrix $A^{(t)}\in\mathbb{Z}^{3\times 3}$ in row-major order:
@@ C.2 Simulating WFAs with RWKV-7
Phase 1 (compute xPxP into scratch). For each $j\in\{1,\dots,n\}$, define $c^{(j)}\in\mathbb{K}^{2n}$ by $c^{(j)}_{i}=P_{i,j}$ for $1\leq i\leq n$ and $c^{(j)}_{i}=0$ for $n+1\leq i\leq 2n$. Let $G_{j}(P):=U(n+j;c^{(j)})$. By Lemma 4, this overwrites scratch coordinate $n+j$ with $\sum_{i=1}^{n}x_{i}P_{i,j}=(xP)_{j}$, leaving all other coordinates unchanged. After $n$ steps, $x\mid s]G_{1}(P)\cdots G_{n}(P)=[x\mid xP]$.
@@ C.2.1 Simulation theorem
Fix $\mathbb{K}=\mathbb{Q}$. Let $\mathcal{A}=(Q,\Sigma,\{M_{\sigma}\},\alpha,\omega)$ be an $n$-state WFA over $\mathbb{K}$ (Definition 13), and let $w=w_{1}\cdots w_{T}$ be an input word. Then there exists a 44-layer RWKV-7 network whose output at position $T$ equals $f_{\mathcal{A}}(w)=\alpha M_{w_{1}}\cdots M_{w_{T}}\omega\in\mathbb{Q}$.
@@
with the convention that empty products are II.
@@ C.3 Iterated $3\times 3$ Matrix Multiplication with RWKV-7
Fix $\mathbb{K}=\mathbb{Q}$. There exists a 44-layer RWKV-7 network such that, given the length-$9N$ stream encoding $A^{(1)},\dots,A^{(N)}$, the network outputs the 99 entries of $P_{N}:=A^{(1)}A^{(2)}\cdots A^{(N)}$ (equivalently, $\operatorname{vec}(P_{N})\in\mathbb{Q}^{9}$) at the final position $t=9N$.
@@
Layers 1–2 compute the position modulo $18$. This yields (i) the within-block offset $\tau\in\{1,\dots,9\}$ and (ii) a block-parity bit $b\in\{0,1\}$ that we use to alternate between two halves of an arithmetic state. Layer 3 stores the last $18$ tokens (two consecutive length-99 blocks) and, from the key $(t\bmod 18,\;w_{t},\dots,w_{t-17})$, look up the RWKV transition parameters needed at time $t$. Concretely, at each token it outputs (in the residual stream) the destination index and coefficient vector for a single dot-product overwrite update implementing one coordinate of the map $x\mapsto x\,B(A^{(\ell-1)})$, where $A^{(\ell-1)}$ is the previous block’s matrix. (We treat the pre-sequence padding as “block 0” encoding $A^{(0)}:=I_{3}$.) Layer 4 maintains an $18$-dimensional row state $x^{(0)}\mid x^{(1)}]\in\mathbb{Q}^{1\times 18}$ inside the first row of a wkv matrix state, where each half has length 99. During each length-99 block, layer 4 applies nine dot-product overwrites that compute a new vector $x_{\text{new}}=x_{\text{old}}\,B(A^{(\ell-1)})$ in the inactive half, using the parity bit $b$ to select which half is source vs. destination. At the final position $t=9N$, the destination half contains $\operatorname{vec}(A^{(1)}\cdots A^{(N-1)})$. A completion readout (computed by the router from the final finite-window key, which contains the entire last block and hence determines $A^{(N)}$) outputs the 99 coordinates of $\operatorname{vec}(A^{(1)}\cdots A^{(N)})$ via nine dot products. ∎
@@
Step 1: reduce $3\times 3$ matrix multiplication to a 99-dimensional linear recurrence.
@@
Right-multiplication by AA acts independently on the three rows of $P$, hence
@@
Thus it suffices to maintain the 99-dimensional state $x_{t}$ and repeatedly apply the linear map $x\mapsto xB(A)$.
@@
Fix a $3\times 3$ matrix AA. The coordinates of $xB(A)$ are length-33 dot products:
@@
We compute these 99 values into the inactive half $x^{(1-b)}$ using 99 overwrite steps. For each pair $(i,j)\in\{1,2,3\}^{2}$, define:
@@
Step 3: align the 99 overwrites with the token stream using a 3-layer router.
@@
Partition the input stream into NN blocks of length 99. For convenience, define an additional “block 0” consisting of padding symbols before the sequence; the router will interpret this as encoding $I_{3}$ and we write $A^{(0)}:=I_{3}$. Let $A^{(\ell)}$ be the matrix encoded by block $\ell$ for $\ell=1,\dots,N$. At any position $t$ in block $\ell$ (so $\ell\in\{1,\dots,N\}$), the model will apply the update corresponding to the previous matrix $A^{(\ell-1)}$ (one-block delay).
@@
Recovering the previous block. At time $t$ in block $\ell$, the last $18$ tokens $(w_{t},w_{t-1},\dots,w_{t-17})$ contain the entire previous block $\ell-1$ (all 99 tokens encoding $A^{(\ell-1)}$), together with a prefix of the current block. Therefore, the tuple
@@
At each position $t$ in block $\ell$, the router provides the overwrite parameters corresponding to the matrix $A^{(\ell-1)}$ and the overwrite index $\tau$. By the discussion above, we choose $A_{t}=U(\mathrm{dst};c)$ so that exactly one coordinate of $p_{t-1}$ is overwritten. After consuming all 99 tokens of block $\ell$, equation (3) implies that the inactive half has been filled with
@@
Since $T$ is the last token of block NN, the router key $(T\bmod 18,\;w_{T},\dots,w_{T-17})$ contains the entire block NN and hence determines $A^{(N)}$. For each $(i,j)\in\{1,2,3\}^{2}$, define the completion vector
@@ C.4 Simulating WFAs with DeltaNet
This subsection shows that a 44-layer stacked DeltaNet can simulate any weighted finite automaton over $\mathbb{Q}$. Throughout this subsection we fix $\mathbb{K}=\mathbb{Q}$ and interpret all arithmetic in $\mathbb{Q}$.
@@ C.4.2 Router primitives: position modulo and token buffering
For any positive integer $m$, there exists a 22-layer stacked DeltaNet whose output includes a one-hot encoding of $\tilde{t}:=t\bmod 2m$ at every position $t$.
@@
The construction mirrors a standard “two-reflections-make-a-rotation” argument. Layer 1 produces the parity bit of $t$ (and can also detect $t=1$ via the BOS token). Layer 2 maintains a 22-dimensional continuous state that advances by a fixed angle every two steps: it alternates between two Householder reflections of the form $I-2\kappa\kappa^{\top}$, which is exactly $H(\beta,k)$ with $\beta=2$ and $k=\kappa$ (and $v_{t}=\mathbf{0}$). The state therefore cycles through $2m$ distinct values, one for each residue class modulo $2m$. A finite MLP on top of a bank of linear readouts can map these $2m$ discrete states to a one-hot encoding of $\tilde{t}$. ∎
@@ C.4.3 Arithmetic primitives realizable by multiplicative DeltaNet steps
Assume $\tfrac{1}{2},\tfrac{1}{3}\in\mathbb{K}$. Fix distinct indices $\mathrm{src},\mathrm{dst},\mathrm{tmp}$ and a scalar $\lambda\in\mathbb{K}$. There exists a length-88 sequence of multiplicative matrices of the form $H(\beta,k)$ such that for every row vector $r$ with $r_{\mathrm{tmp}}=0$, the resulting row vector $r^{\prime}$ satisfies:
@@
    Copy $\mathrm{src}$ into $\mathrm{tmp}$ by a unit transvection $T(\mathrm{src}\to\mathrm{tmp})$, implemented using 33 steps via Lemma 10.
@@
    Add $\mathrm{tmp}$ into $\mathrm{dst}$ by a unit transvection $T(\mathrm{tmp}\to\mathrm{dst})$, implemented using 33 steps via Lemma 10.
@@
Phase 2 (accumulate xPxP into scratch).
@@
Each scaled add costs 88 multiplicative steps and leaves temp equal to 0. After all $n^{2}$ updates, scratch equals xPxP while main remains $x$.
@@
For each $j=1,\dots,n$, add scratch coordinate $n+j$ into main coordinate $j$ using a unit transvection $T(n+j\to j)$, implemented by 33 multiplicative steps via Lemma 10. Since main is 0, after these $3n$ steps we have main =xP=xP and scratch remains xPxP.
@@
C.4.4 Simulation theorem: a 44-layer DeltaNet simulates any WFA

### 05 Mar 2026 v2

@@ 2.3 Linear RNNs
Definition 4 leaves open the specific parameterization of $A$ and $b$ as a function of $\mathbf{x}_{t}$. The choice of parameterization is crucial for theoretical expressivity (Merrill et al., 2024) as well practical concerns like memory. For instance, constrained parameterizations (time-invariant or diagonal) were explored in early LRNNs, but these were shown to limit expressivity to $\mathsf{TC}^{0}$ (Merrill et al., 2024). Since then, newer LRNNs have adopted richer parameterizations that are non-diagonal but still efficient. Two broad classes of “slightly non-diagonal” parameterizations are diagonal-plus-low-rank (DPLR) and permutation-diagonal (PD) LRNNS. [^2] Henceforth, for brevity, we will write $\mathbf{A}_{t}$, $\mathbf{k}_{t}$, etc. to mean $A_{t}(\mathbf{x}_{t})$, $k_{t}(\mathbf{x}_{t})$, etc., leaving the dependence on $\mathbf{x}_{t}$ implicit.
@@ 5.1 DPLR LRNNs are $\mathsf{PNC}^{1}$-Complete
The input is a token stream of length $9n$ over $\{-1,0,1\}$, partitioned into $n$ consecutive blocks of length $9$. Block $t$ encodes a matrix $\mathbf{A}_{t}\in\mathbb{K}^{3\times 3}$. The iterated matrix multiplication problem is to compute $\mathbf{P}_{n}=\prod_{t=1}^{n}\mathbf{A}_{t}\in\mathbb{Z}^{3\times 3}$.
@@
There exist a $4$-layer RWKV-7 and a $4$-layer DeltaNet that solves the iterated $3\times 3$ matrix multiplication problem and, in particular, checks if the $(0,0)$ entry of $\mathbf{P}_{n}$ is positive.
@@
For any $n$-state WFA $\mathcal{A}$ over $\mathbb{Q}$, there exist a $4$-layer RWKV-7 and a 4-layer DeltaNet that both compute $f_{\mathcal{A}}(w)$.
@@ 6 Experiments
To better evaluate learning capacity, we construct balanced datasets with length generalization. The train set contains 70K samples uniformly drawn from $N\in[1,100]$ and the validation set contains 20K samples drawn from the same range. We evaluated on three test splits, each containing 10K samples, drawn from $[1,100]$, $[101,200]$, and $[201,300]$. Here, $N$ denotes the number of nodes in the graph task and the size of the matrix in the matrix multiplication task.
@@ A.2 Logarithmic Precision
The above reduction can easily be seen to be implementable by a logspace-uniform $\mathsf{AC}^{0}$ reduction. The values $hn$ (for each $h=1,\ldots,m$) can be precomputed by a logspace machine since this depends only on the input size (more precisely, one can enumerate enough of $hn$ for up to $m=O(n)$). The $\mathsf{AC}^{0}$ circuit simply needs to perform simple additions. ∎
@@
At the following separator it tests whether $S=0$ using the zero-mask. If $S=0$ (i.e. $s_{m-1}=i_{m}$), it moves to a state $q_{j}^{=}$ and on each symbol of the $j_{m}$-block simply does $S\leftarrow S+1$ (leaving $I$ unchanged); since $S$ starts at 0, after $0^{j_{m}}$ we have $S=j_{m}$, and at the next separator it resets $I\leftarrow 0$ and returns to $q_{\mathsf{edge}}$, so $s_{m}=j_{m}$. If instead $S\neq 0$, the machine moves to a state $q_{j}^{\neq}$. On each symbol of the $j_{m}$-block it now performs
@@
where the choice again uses only the zero-mask bit of $I$. Initially $I=i_{m}$, so after $\min\{i_{m},j_{m}\}$ symbols we have added that many units back to $S$ and subtracted them from $I$. Because $j_{m}\geq i_{m}$, by the end of the $j_{m}$-block we have $I=0$ and
@@ C.1 Preliminaries
The input is a token stream of length $9N$ over alphabet $\mathbb{Z}$, partitioned into $N$ consecutive blocks of length $9$. The $t$-th block encodes a matrix $A^{(t)}\in\mathbb{Z}^{3\times 3}$ in row-major order:
@@ C.2 Simulating WFAs with RWKV-7
Phase 1 (compute $xP$ into scratch). For each $j\in\{1,\dots,n\}$, define $c^{(j)}\in\mathbb{K}^{2n}$ by $c^{(j)}_{i}=P_{i,j}$ for $1\leq i\leq n$ and $c^{(j)}_{i}=0$ for $n+1\leq i\leq 2n$. Let $G_{j}(P):=U(n+j;c^{(j)})$. By Lemma 4, this overwrites scratch coordinate $n+j$ with $\sum_{i=1}^{n}x_{i}P_{i,j}=(xP)_{j}$, leaving all other coordinates unchanged. After $n$ steps, $x\mid s]G_{1}(P)\cdots G_{n}(P)=[x\mid xP]$.
@@ C.2.1 Simulation theorem
Fix $\mathbb{K}=\mathbb{Q}$. Let $\mathcal{A}=(Q,\Sigma,\{M_{\sigma}\},\alpha,\omega)$ be an $n$-state WFA over $\mathbb{K}$ (Definition 13), and let $w=w_{1}\cdots w_{T}$ be an input word. Then there exists a $4$-layer RWKV-7 network whose output at position $T$ equals $f_{\mathcal{A}}(w)=\alpha M_{w_{1}}\cdots M_{w_{T}}\omega\in\mathbb{Q}$.
@@
with the convention that empty products are $I$.
@@ C.3 Iterated $3\times 3$ Matrix Multiplication with RWKV-7
Fix $\mathbb{K}=\mathbb{Q}$. There exists a $4$-layer RWKV-7 network such that, given the length-$9N$ stream encoding $A^{(1)},\dots,A^{(N)}$, the network outputs the $9$ entries of $P_{N}:=A^{(1)}A^{(2)}\cdots A^{(N)}$ (equivalently, $\operatorname{vec}(P_{N})\in\mathbb{Q}^{9}$) at the final position $t=9N$.
@@
Layers 1–2 compute the position modulo $18$. This yields (i) the within-block offset $\tau\in\{1,\dots,9\}$ and (ii) a block-parity bit $b\in\{0,1\}$ that we use to alternate between two halves of an arithmetic state. Layer 3 stores the last $18$ tokens (two consecutive length-$9$ blocks) and, from the key $(t\bmod 18,\;w_{t},\dots,w_{t-17})$, look up the RWKV transition parameters needed at time $t$. Concretely, at each token it outputs (in the residual stream) the destination index and coefficient vector for a single dot-product overwrite update implementing one coordinate of the map $x\mapsto x\,B(A^{(\ell-1)})$, where $A^{(\ell-1)}$ is the previous block’s matrix. (We treat the pre-sequence padding as “block 0” encoding $A^{(0)}:=I_{3}$.) Layer 4 maintains an $18$-dimensional row state $x^{(0)}\mid x^{(1)}]\in\mathbb{Q}^{1\times 18}$ inside the first row of a wkv matrix state, where each half has length $9$. During each length-$9$ block, layer 4 applies nine dot-product overwrites that compute a new vector $x_{\text{new}}=x_{\text{old}}\,B(A^{(\ell-1)})$ in the inactive half, using the parity bit $b$ to select which half is source vs. destination. At the final position $t=9N$, the destination half contains $\operatorname{vec}(A^{(1)}\cdots A^{(N-1)})$. A completion readout (computed by the router from the final finite-window key, which contains the entire last block and hence determines $A^{(N)}$) outputs the $9$ coordinates of $\operatorname{vec}(A^{(1)}\cdots A^{(N)})$ via nine dot products. ∎
@@
Step 1: reduce $3\times 3$ matrix multiplication to a $9$-dimensional linear recurrence.
@@
Right-multiplication by $A$ acts independently on the three rows of $P$, hence
@@
Thus it suffices to maintain the $9$-dimensional state $x_{t}$ and repeatedly apply the linear map $x\mapsto xB(A)$.
@@
Fix a $3\times 3$ matrix $A$. The coordinates of $xB(A)$ are length-$3$ dot products:
@@
We compute these $9$ values into the inactive half $x^{(1-b)}$ using $9$ overwrite steps. For each pair $(i,j)\in\{1,2,3\}^{2}$, define:
@@
Step 3: align the $9$ overwrites with the token stream using a 3-layer router.
@@
Partition the input stream into $N$ blocks of length $9$. For convenience, define an additional “block 0” consisting of padding symbols before the sequence; the router will interpret this as encoding $I_{3}$ and we write $A^{(0)}:=I_{3}$. Let $A^{(\ell)}$ be the matrix encoded by block $\ell$ for $\ell=1,\dots,N$. At any position $t$ in block $\ell$ (so $\ell\in\{1,\dots,N\}$), the model will apply the update corresponding to the previous matrix $A^{(\ell-1)}$ (one-block delay).
@@
Recovering the previous block. At time $t$ in block $\ell$, the last $18$ tokens $(w_{t},w_{t-1},\dots,w_{t-17})$ contain the entire previous block $\ell-1$ (all $9$ tokens encoding $A^{(\ell-1)}$), together with a prefix of the current block. Therefore, the tuple
@@
At each position $t$ in block $\ell$, the router provides the overwrite parameters corresponding to the matrix $A^{(\ell-1)}$ and the overwrite index $\tau$. By the discussion above, we choose $A_{t}=U(\mathrm{dst};c)$ so that exactly one coordinate of $p_{t-1}$ is overwritten. After consuming all $9$ tokens of block $\ell$, equation (3) implies that the inactive half has been filled with
@@
Since $T$ is the last token of block $N$, the router key $(T\bmod 18,\;w_{T},\dots,w_{T-17})$ contains the entire block $N$ and hence determines $A^{(N)}$. For each $(i,j)\in\{1,2,3\}^{2}$, define the completion vector
@@ C.4 Simulating WFAs with DeltaNet
This subsection shows that a $4$-layer stacked DeltaNet can simulate any weighted finite automaton over $\mathbb{Q}$. Throughout this subsection we fix $\mathbb{K}=\mathbb{Q}$ and interpret all arithmetic in $\mathbb{Q}$.
@@ C.4.2 Router primitives: position modulo and token buffering
For any positive integer $m$, there exists a $2$-layer stacked DeltaNet whose output includes a one-hot encoding of $\tilde{t}:=t\bmod 2m$ at every position $t$.
@@
The construction mirrors a standard “two-reflections-make-a-rotation” argument. Layer 1 produces the parity bit of $t$ (and can also detect $t=1$ via the BOS token). Layer 2 maintains a $2$-dimensional continuous state that advances by a fixed angle every two steps: it alternates between two Householder reflections of the form $I-2\kappa\kappa^{\top}$, which is exactly $H(\beta,k)$ with $\beta=2$ and $k=\kappa$ (and $v_{t}=\mathbf{0}$). The state therefore cycles through $2m$ distinct values, one for each residue class modulo $2m$. A finite MLP on top of a bank of linear readouts can map these $2m$ discrete states to a one-hot encoding of $\tilde{t}$. ∎
@@ C.4.3 Arithmetic primitives realizable by multiplicative DeltaNet steps
Assume $\tfrac{1}{2},\tfrac{1}{3}\in\mathbb{K}$. Fix distinct indices $\mathrm{src},\mathrm{dst},\mathrm{tmp}$ and a scalar $\lambda\in\mathbb{K}$. There exists a length-$8$ sequence of multiplicative matrices of the form $H(\beta,k)$ such that for every row vector $r$ with $r_{\mathrm{tmp}}=0$, the resulting row vector $r^{\prime}$ satisfies:
@@
    Copy $\mathrm{src}$ into $\mathrm{tmp}$ by a unit transvection $T(\mathrm{src}\to\mathrm{tmp})$, implemented using $3$ steps via Lemma 10.
@@
    Add $\mathrm{tmp}$ into $\mathrm{dst}$ by a unit transvection $T(\mathrm{tmp}\to\mathrm{dst})$, implemented using $3$ steps via Lemma 10.
@@
Phase 2 (accumulate $xP$ into scratch).
@@
Each scaled add costs $8$ multiplicative steps and leaves temp equal to 0. After all $n^{2}$ updates, scratch equals $xP$ while main remains $x$.
@@
For each $j=1,\dots,n$, add scratch coordinate $n+j$ into main coordinate $j$ using a unit transvection $T(n+j\to j)$, implemented by $3$ multiplicative steps via Lemma 10. Since main is 0, after these $3n$ steps we have main =xP=xP and scratch remains $xP$.
@@
C.4.4 Simulation theorem: a $4$-layer DeltaNet simulates any WFA